# Pierre Auger Collaboration

### v1

@@ I Introduction
Due to the rapid decrease in particle flux at ultra-high energies, modern cosmic-ray observatories perform indirect measurements of the rare particles by detecting generated air showers instead. The influence of the primary mass on the shower development can be characterized mainly by the number of muons and the atmospheric depth of the shower maximum $X_{\mathrm{max}}$ at which the shower reaches its maximum size. With increasing primary mass, the number of induced sub-showers increases, and the energy per nucleon reduces, leading to an $X_{\mathrm{max}}$ higher up in the atmosphere and decreasing shower fluctuations. The increase in the number of sub-showers additionally causes an increase in the number of produced muons. Since the current generation of hadronic interaction models cannot describe the muon component in full detail [11, 12, 13, 14], currently the most precise composition studies at ultra-high energies rely on measurements of $X_{\mathrm{max}}$.
@@
With the advent of deep learning, new possibilities have emerged for designing learning algorithms, i.e., deep neural networks (DNNs), to analyze high-dimensional and complex data in computational sciences [23, 24] as well as in physics [25]. Trained on large simulation libraries, these algorithms are capable of recognizing small patterns to which conventional methods were previously not sensitive. This recent progress provided improved reconstruction algorithms in astroparticle physics, e.g., imaging air Cherenkov telescopes [26], gravitational wave detection [27], neutrino [28, 29, 30] and cosmic-ray observatories [31], including the reconstruction of $X_{\mathrm{max}}$ [32, 33] and other air shower properties [34, 35]. So far, the potential of deep-learning-based methods for improved reconstruction in astroparticle physics has been demonstrated, but the application to measured data, including a comprehensive study of systematic uncertainties and associated new insights, is limited. This work aims to close this gap and shows a successful application to measured data starting at the raw detector signals.
@@ II.1 Surface Detector data
Figure 2: Simulated signal traces of a cosmic-ray event in the SD stations before (left) and after the pre-processing (right). (a) Simulated signal trace of a station close to the shower core (top), at a distance of around 1000 m (middle), and at a distance of around 3500 m to the shower core (bottom). Different colors indicate signals from different shower components. The black line denotes the total measured signal, including the saturation effects of the electronics that are only simulated for the sum of all shower components.
@@ III.1 Deep-learning-based reconstruction
The DNN trained for the reconstruction of $X_{\mathrm{max}}$ is based on the signal traces measured using the WCDs of the triangular grid of the SD array and the arrival times. To process the temporal and spatial structure of the particle footprint, the DNN uses the following architecture methodology, separating the analysis in space and time. Since the SD grid is triangular with a regular spacing of 1500 m, we use the axial representation for re-indexing into a Cartesian grid [31]. For a memory-efficient re-indexing, we use a cutout of $13\times 13$ stations, where the station with the largest signal defines the center of this grid. The dimensions of $13\times 13$ stations guarantee that, on average, more than $99.9\%$ of the triggered stations per event are contained within this sub-array. See Fig. 1(a) that visualizes this process using an example SD event.
@@
To analyze the temporal structure of the signal traces, recurrent long short-term memory (LSTM) layers [42] are utilized in the first part of the network. The identical network subpart with the same adaptive parameters is applied to each signal trace, i.e., we apply weight sharing along all stations as similar particles induce similar responses in the detector. The output of this network can be interpreted as an image of $13\times 13$ pixels (stations) with ten channels instead of three in a natural RGB image, as the recurrent network part characterizes the traces of each station into ten features. These features are the input for the convolutional part and are concatenated with the channel of arrival times and an additional channel characterizing the detector states of the surrounding WCDs (working/not working). The next stage is based on an advanced type of convolutional neural network (CNN) [23] to exploit the spatial structure of the event. We make use of the so-called HexaConv layers [43] and residual connections [44, 45], which extends the principle of a filter sliding along an image by a rotation. This is a meaningful extension as the induced signal patterns are to first order independent of the azimuth angle. Finally, after a ResNet-like architecture, a single node for the prediction of $X_{\mathrm{max}}$ forms the output of the DNN. For a detailed description of the DNN architecture, we refer to Ref. [32].
@@ III.3 Reconstruction of the $\mathbf{\textit{X}_{\text{max}}}$ moments
Thus, to reconstruct $\sigma(X_{\mathrm{max}})$, not only the resolution but due to the composition dependence, i.e., $X_{\mathrm{max}}$ dependence also, the covariance term has to be considered. In the absence of sufficient information, estimators like DNNs trained with the mean-squared-error objective function tend to predict samples close to the mean since the reconstruction is ambiguous. Therefore, at low energies in particular, the DNN is likely to reconstruct, on average, iron with positive and proton with negative $X_{\mathrm{max}}$ bias. In turn, the covariance term is negative for the DNN, i.e., it will in part cancel the resolution term. We studied the dependence of the sum of both terms, the covariance and the resolution for various hadronic models and compositions as a function of energy. In the case of our trained DNN, we found that, to a good approximation, both terms cancel or are small in comparison to the physical fluctuations in $X_{\mathrm{max}}$. Therefore, the standard deviation of the distribution formed by the DNN predictions is used as the estimate for $\sigma(X_{\mathrm{max}})$, and the statistical uncertainty on $\sigma(X_{\mathrm{max}})$ is obtained using bootstrapping. Deviations, i.e., scenarios with non-canceling contributions of the resolution and covariance, will translate into a composition bias of the second moment in this forward-folding approach and propagate into the systematic uncertainties of the $\sigma(X_{\mathrm{max}})$ measurement.
@@
In Fig. 4, the performance in reconstructing the evolution of the moments $\langle X_{\mathrm{max}}\rangle$ (top) and $\sigma(X_{\mathrm{max}})$ (bottom) using the SD is depicted for three different scenarios. A pure proton composition is shown in red, pure iron in blue, and the Auger mix in yellow for the three hadronic interaction models EPOS-LHC (filled diamonds), QGSJetII-04 (grey-filled squares), and Sibyll2.3c (open circles). As a reference, the injected (true) moments are shown as white boxes where their vertical sizes indicate the statistical uncertainty prior to the $X_{\mathrm{max}}$ reconstruction by the DNN and the energy reconstruction [38]. Note that only EPOS-LHC was used as a hadronic interaction model for training the DNN. Since a large fraction of the EPOS-LHC simulations was used for the DNN training, the statistical uncertainty is larger for EPOS-LHC events than for QGSJetII-04 and Sibyll2.3c, where all simulations could be used for testing. We find that the performance in the determination of $\langle X_{\mathrm{max}}\rangle$ depends on energy and the hadronic interaction models. Above 10 EeV the performance differences across the models and mass composition scenarios are small  [^6] A residual plot for the three investigated models can be found in Fig. 17. The reconstruction bias shows a dependence on the composition at low energies. At $3$ EeV, the pure iron composition is subject to a bias of up to 10to $1510\;\;\mathrm{to}\;$ 15 $\mathrm{g\,cm^{-2}}$ that reduces with energy. For a pure proton composition, a similar dependence is visible, of up to $-10~{}\mathrm{g\,cm^{-2}}$ decreasing with energy.
For the Auger mix, the composition bias is independent of energy for QGSJetII-04 and Sibyll2.3c. Only for EPOS-LHC, a small composition bias of up to $-5$ $\mathrm{g\,cm^{-2}}$ can be seen below $6$ EeV. Otherwise, we find an interaction-model bias of $-5$ $\mathrm{g\,cm^{-2}}$  for QGSJetII-04 that applies to all compositions. For Sibyll2.3c, this bias amounts to $-12$ $\mathrm{g\,cm^{-2}}$. Because EPOS-LHC was used for training, no such bias is visible for this model. Both the composition bias and the hadronic-interaction model bias would propagate into the systematic uncertainty of the $\langle X_{\mathrm{max}}\rangle$ measurement. However, since the Observatory features a hybrid detector design, both biases, and their energy dependence can be removed (as investigated in a simulation study) by re-calibrating the DNN measurement using fluorescence observations (as discussed in Section III.4.2).
@@
The composition bias of the $\sigma(X_{\mathrm{max}})$ reconstruction as seen in Fig. 4 (bottom) depends on energy and is below $5~{}\mathrm{g\,cm^{-2}}$ above 10 EeV for proton and iron. For the Auger mix it is even lower at these high energies and not significant. Overall, the biases observed for reconstructing $\sigma(X_{\mathrm{max}})$ are small over a large range of energies. For this reason, no calibration using the FD will be performed. Furthermore, the biases found here will be transferred to the systematic uncertainty of our composition measurements. Different from the measurement of $\langle X_{\mathrm{max}}\rangle$, the estimation of $\sigma(X_{\mathrm{max}})$ is not subject to a strong dependence on the hadronic interaction model. This can be explained by the fact that a large part of the shower fluctuations depends on the fluctuations of $X_{1}$, i.e., the traversed depth prior to the first interaction, owing to the given nuclear cross sections with air molecules. Depending on the primary, the expected fluctuations in $X_{1}$ are in the order of 10 $\mathrm{g\,cm^{-2}}$ (50 $\mathrm{g\,cm^{-2}}$) for iron (proton) nuclei.
@@ III.4 Calibration
where $f_{\Delta X_{\mathrm{max}}}(y)$ denotes the dependence of the $X_{\mathrm{max}}$ prediction on the variable $y$. This approach performed for each event separately ensures meaningful corrections of the predictions beyond the first moment of $X_{\mathrm{max}}$.
Finally, using hybrid events, we calibrate the $X_{\mathrm{max}}$ predictions to the scale of the FD and remove the dependence on the hadronic interaction model used during the algorithm training, the composition, and any remaining differences in the detector simulation.
@@ III.4.1 Corrections using Surface Detector data
The WCD stations of the SD are exposed to a harsh environment, with changes in temperature covering several tens of ${}^{\circ}\mathrm{C}$. During the many years of operation, the PMTs, read-out electronics, water, and reflective liner are subject to aging effects. By utilizing muons that constantly cross the detector stations, the average shape of the single-muon signal is monitored using the area over peak ratio $A/P$, which relates the deposited charge in the detector (integrated pulse) to its height and is a rough measure of the signal duration. Since $A/P$ is a monitoring observable summarizing the characteristics of the individual PMT responses, water quality, and liner reflectivity, it is specific to every station and changes as a function of time.
@@
When monitoring the distribution of $A/P$ values of all stations, over the years, a decrease in its average $\langle A/P\rangle$ from $3.20$ to $2.95$ can be observed [55, 56]. We assume this to be mainly caused by the decrease in the liner reflectivity or water transparency, leading to a drift of the $X_{\mathrm{max}}$ predictions as a function of time (see Fig. 5(a)). Since our simulation library is currently limited to simulated stations with $A/P=3.2$, the predictions of the DNN have to be calibrated as a function of the $\langle A/P\rangle$ — the average $A/P$ of all triggered SD stations in a given event — to remove possible time dependencies of the predictions. As depicted in Fig. 5(c), the dependence of the $X_{\mathrm{max}}$ predictions on $\langle A/P\rangle$ can be modeled linearly and does not depend on energy. Additionally, we find an increase of this dependence as a function of the zenith angle, likely caused by the fact that the average distance a particle travels through the detector rises with the zenith angle. By parameterizing (blue line in Fig. 5(d)) this dependence as $c^{*}(\theta)=\alpha\,\cos(\theta)+\beta$, with $\alpha=(-135.1\pm 9.2)$ $\mathrm{g\,cm^{-2}}$, and $\beta=(152.6\pm 7.2)$ $\mathrm{g\,cm^{-2}}$, we calibrate the predictions using:
@@
Figure 8: Application of the DNN $X_{\mathrm{max}}$ estimation to hybrid data. (a) Correlation between FD observations and DNN predictions using SD data. (b) The energy-dependent bias of the SD-based reconstruction of the DNN when compared to the reconstruction of the fluorescence detector.
@@
The change in pressure and temperature due to diurnal and seasonal variations causes small influences on the detector response and the conversion of distance to $X_{\mathrm{max}}$, hence affecting its reconstruction. To remove seasonal and diurnal variations, we investigated the reconstruction as a function of time on a yearly and daily basis [^7] The time is given in UTC., as shown in Fig. 7(a) and Fig. 7(b). We find small variations of the size of $2~{}\mathrm{g\,cm^{-2}}$ or $1~{}\mathrm{g\,cm^{-2}}$, respectively. We first calibrate the $X_{\mathrm{max}}$ predictions to remove the seasonal variation by fitting a sine function to the data and correcting it by using the time of year on an event-by-event basis. These predictions are also used to remove the diurnal variations by again fitting a sine wave and correcting events using Eq. 4. The dependencies after correction are depicted in Fig. 7(d) and Fig. 7(e).

### v2

@@ I Introduction
Due to the rapid decrease in particle flux at ultra-high energies, modern cosmic-ray observatories perform indirect measurements of the rare particles by detecting generated air showers instead. The influence of the primary mass on the shower development can be characterized mainly by the number of muons and the atmospheric depth of the shower maximum $X_{\mathrm{max}}$ at which the shower reaches its maximum size. At a given primary energy, with increasing primary mass, the number of induced sub-showers increases, and the energy per nucleon reduces, leading to an $X_{\mathrm{max}}$ higher up in the atmosphere and decreasing shower fluctuations. The increase in the number of sub-showers additionally causes an increase in the number of produced muons. Since the current generation of hadronic interaction models cannot describe the muon component in full detail [11, 12, 13, 14], currently the most precise composition studies at ultra-high energies rely on measurements of $X_{\mathrm{max}}$.
@@
With the advent of deep learning, new possibilities have emerged for designing learning algorithms, i.e., deep neural networks (DNNs), to analyze high-dimensional and complex data in computational sciences [23, 24] as well as in physics [25]. Trained on large simulation libraries, these algorithms are capable of recognizing small patterns [26], like complex mesh structures, skin texture, or dog faces, to which conventional methods were previously not sensitive. This recent progress provided improved reconstruction algorithms in astroparticle physics, e.g., imaging air Cherenkov telescopes [27], gravitational wave detection [28], neutrino [29, 30, 31] and cosmic-ray observatories [32], including the reconstruction of $X_{\mathrm{max}}$ [33, 34] and other air shower properties [35, 36]. So far, the potential of deep-learning-based methods for improved reconstruction in astroparticle physics has been demonstrated, but the application to measured data, including a comprehensive study of systematic uncertainties and associated new insights, is limited. This work aims to close this gap and shows a successful application to measured data starting at the raw detector signals.
@@ II.1 Surface Detector data
Figure 2: Simulated signal traces of a cosmic-ray event in the SD stations before (left) and after the pre-processing (right). Simulated signal trace of a station close to the shower core (top), at a distance of around 1000 m (middle), and at a distance of around 3500 m to the shower core (bottom). Different colors indicate signals from different shower components. The black line denotes the total measured signal, including the saturation effects of the electronics that are only simulated for the sum of all shower components.
@@ II.2.2 Hybrid measurements
Figure 3: DNN Architecture used to reconstruct $X_{\mathrm{max}}$. The numbers in brackets denote the output shapes.
@@ III.1 Deep-learning-based reconstruction
The DNN trained for the reconstruction of $X_{\mathrm{max}}$ is based on the signal traces measured using the WCDs of the triangular grid of the SD array and the arrival times. To process the temporal and spatial structure of the particle footprint, the DNN uses the following architecture methodology, shown in Fig. 3, separating the analysis in space and time. Since the SD grid is triangular with a regular spacing of 1500 m, we use the axial representation for re-indexing into a Cartesian grid [32]. For a memory-efficient re-indexing, we use a cutout of $13\times 13$ stations, where the station with the largest signal defines the center of this grid. The dimensions of $13\times 13$ stations guarantee that, on average, more than $99.9\%$ of the triggered stations per event are contained within this sub-array. See Fig. 1(a) that visualizes this process using an example SD event.
@@
To characterize the temporal structure of the signal traces, recurrent long short-term memory (LSTM) layers [43] are utilized in the first part of the network. The identical network subpart with the same adaptive parameters is applied to each signal trace, i.e., we apply weight sharing along all stations as similar particles induce similar responses in the detector. The output of this network can be interpreted as an image of $13\times 13$ pixels (stations) with ten channels instead of three in a natural RGB image, as the recurrent network part characterizes the traces of each station into ten features. These features are the input for the convolutional part and are concatenated with the channel of arrival times and an additional channel characterizing the detector states of the surrounding WCDs (working/not working). The next stage is based on an advanced type of convolutional neural network (CNN) [23] to exploit the spatial structure of the event. We make use of the so-called HexaConv layers [44] and residual connections [45, 46], which extends the principle of a filter sliding along an image by a rotation. This is a meaningful extension as the induced signal patterns are to first order independent of the azimuth angle. Finally, after a ResNet-like architecture, a single node for the prediction of $X_{\mathrm{max}}$ forms the output of the DNN. For a detailed description of the DNN architecture, we refer to Ref. [33].
@@ III.3 Reconstruction of the $\mathbf{\textit{X}_{\text{max}}}$ moments
To reconstruct $\sigma(X_{\mathrm{max}})$, the resolution and the covariance term have to be considered since the reconstruction shows a composition, i.e., $X_{\mathrm{max}}$ dependence. In the absence of sufficient information, estimators like DNNs trained with the mean-squared-error objective function tend to predict samples close to the mean since the reconstruction is ambiguous. Therefore, particularly at low energies, the DNN is likely to reconstruct, on average, iron with positive and proton with negative $X_{\mathrm{max}}$ bias. In turn, the covariance term is negative for the DNN, i.e., it will in part cancel the resolution term. We studied the dependence of the sum of both terms, the covariance and the resolution for various hadronic models and compositions as a function of energy. In the case of our trained DNN, we found that, to a good approximation, both terms cancel or are small in comparison to the physical fluctuations in $X_{\mathrm{max}}$. Therefore, the standard deviation of the distribution formed by the DNN predictions is used as the estimate for $\sigma(X_{\mathrm{max}})$, and the statistical uncertainty on $\sigma(X_{\mathrm{max}})$ is obtained using bootstrapping. Deviations, i.e., scenarios with non-canceling contributions of the resolution and covariance, will translate into a composition bias of the second moment in this forward-folding approach and propagate into the systematic uncertainties of the $\sigma(X_{\mathrm{max}})$ measurement.
@@
In Fig. 5, the performance in reconstructing the evolution of the moments $\langle X_{\mathrm{max}}\rangle$ (top) and $\sigma(X_{\mathrm{max}})$ (bottom) using the SD is depicted for three different scenarios. A pure proton composition is shown in red, pure iron in blue, and the Auger mix in yellow for the three hadronic interaction models EPOS-LHC (filled diamonds), QGSJetII-04 (grey-filled squares), and Sibyll2.3c (open circles). As a reference, the injected (true) moments are shown as white boxes where their vertical sizes indicate the statistical uncertainty prior to the $X_{\mathrm{max}}$ reconstruction by the DNN and the energy reconstruction [39]. Note that only EPOS-LHC was used as a hadronic interaction model for training the DNN. Since a large fraction of the EPOS-LHC simulations was used for the DNN training, the statistical uncertainty is larger for EPOS-LHC events than for QGSJetII-04 and Sibyll2.3c, where all simulations could be used for testing. We find that the performance in the determination of $\langle X_{\mathrm{max}}\rangle$ depends on energy and the hadronic interaction models. An interaction-model bias, i.e., a systematic shift for all compositions, of $-5$ $\mathrm{g\,cm^{-2}}$ is visible for QGSJetII-04. For Sibyll2.3c, this bias amounts to $-12$ $\mathrm{g\,cm^{-2}}$. Because EPOS-LHC was used for training, no such bias is visible for this model. Above 10 EeV the performance differences across the models and mass composition scenarios are small  [^6] A residual plot for the three investigated models can be found in Fig. 18. The reconstruction bias shows a dependence on the composition at low energies. At $3$ EeV, the pure iron composition is subject to a bias of up to 10to $1510\;\;\mathrm{to}\;$ 15 $\mathrm{g\,cm^{-2}}$ that reduces with energy. For a pure proton composition, a similar dependence is visible, of up to $-10~{}\mathrm{g\,cm^{-2}}$ decreasing with energy.
For the Auger mix, the composition bias is independent of energy for QGSJetII-04 and Sibyll2.3c. Only for EPOS-LHC, a small composition bias of up to $-5$ $\mathrm{g\,cm^{-2}}$ can be seen below $6$ EeV. Both the composition bias and the hadronic-interaction model bias would propagate into the systematic uncertainty of the $\langle X_{\mathrm{max}}\rangle$ measurement. However, since the Observatory features a hybrid detector design, both biases, and their energy dependence can be removed (as investigated in a simulation study) by re-calibrating the DNN measurement using fluorescence observations (as discussed in Section III.4.2).
@@
The composition bias of the $\sigma(X_{\mathrm{max}})$ reconstruction as seen in Fig. 5 (bottom) depends on energy and is below $5~{}\mathrm{g\,cm^{-2}}$ above 10 EeV for proton and iron. Only for iron in the QGSJetII-04 model, a bias around 5 $\mathrm{g\,cm^{-2}}$ can be seen above 10 EeV. For the Auger mix it is even lower at these high energies and not significant. Overall, the biases observed for reconstructing $\sigma(X_{\mathrm{max}})$ are small over a large range of energies. For this reason, no calibration using the FD will be performed. Furthermore, the biases found here will be transferred to the systematic uncertainty of our composition measurements. Different from the measurement of $\langle X_{\mathrm{max}}\rangle$, the estimation of $\sigma(X_{\mathrm{max}})$ is not subject to a strong dependence on the hadronic interaction model. This can be explained by the fact that a large part of the shower fluctuations depends on the fluctuations of $X_{1}$, i.e., the traversed depth prior to the first interaction, owing to the given nuclear cross sections with air molecules. Depending on the primary, the expected fluctuations in $X_{1}$ are in the order of 10 $\mathrm{g\,cm^{-2}}$ (50 $\mathrm{g\,cm^{-2}}$) for iron (proton) nuclei.
@@ III.4 Calibration
where $f_{\Delta X_{\mathrm{max}}}(y)$ denotes the dependence of the $X_{\mathrm{max}}$ prediction on the variable $y$. This approach performed for each event separately ensures meaningful corrections of the predictions beyond the first moment of $X_{\mathrm{max}}$. Finally, using hybrid events, we calibrate the $X_{\mathrm{max}}$ predictions to the scale of the FD and remove the dependence on the hadronic interaction model used during the algorithm training, the composition, and any remaining differences between the measured data and the detector simulation.
@@ III.4.1 Corrections using Surface Detector data
The WCD stations of the SD are exposed to a harsh environment, with changes in temperature covering several tens of ${}^{\circ}\mathrm{C}$. During the many years of operation, the PMTs, read-out electronics, water, and reflective liner are subject to aging effects. By utilizing muons that constantly cross the detector stations, the average shape of the single-muon signal is monitored using the area over peak ratio $A/P$, which relates the deposited charge in the detector (integrated pulse) to its height and is a rough measure of the signal duration. Since $A/P$ is a monitoring observable summarizing the characteristics of the individual PMT responses, water quality, and liner reflectivity, it is specific to every station and changes with time.
@@
When monitoring the distribution of $A/P$ values of all stations, over the years, a decrease in its average $\langle A/P\rangle$ from $3.20$ to $2.95$ can be observed [56, 57]. We assume this to be mainly caused by the decrease in the liner reflectivity or water transparency, leading to a drift of the $X_{\mathrm{max}}$ predictions as a function of time (see Fig. 6(a)). Since our simulation library is currently limited to simulated stations with $A/P=3.2$, the predictions of the DNN have to be calibrated as a function of the $\langle A/P\rangle$ — the average $A/P$ of all triggered SD stations in a given event — to remove possible time dependencies of the predictions. As depicted in Fig. 6(c), the dependence of the $X_{\mathrm{max}}$ predictions on $\langle A/P\rangle$ can be modeled linearly and was found to not depend on energy. Additionally, we find an increase of this dependence as a function of the zenith angle, likely caused by the fact that the average distance a particle travels through the detector rises with the zenith angle. By parameterizing (blue line in Fig. 6(d)) this dependence as $c^{*}(\theta)=\alpha\,\cos(\theta)+\beta$, with $\alpha=(-135.1\pm 9.2)$ $\mathrm{g\,cm^{-2}}$, and $\beta=(152.6\pm 7.2)$ $\mathrm{g\,cm^{-2}}$, we calibrate the predictions using:
@@
Figure 9: Application of the DNN $X_{\mathrm{max}}$ estimation to hybrid data. (a) Correlation between FD observations and DNN predictions using SD data. Furthermore, bias $\mu$, resolution $\sigma$, Pearson correlation coefficient $\rho$, and the number of events $N$ are given. (b) The energy-dependent bias of the SD-based reconstruction of the DNN when compared to the reconstruction of the FD.
@@
The change in pressure and temperature due to diurnal and seasonal variations causes small influences on the detector response and the conversion of distance to $X_{\mathrm{max}}$, hence affecting its reconstruction. To remove seasonal and diurnal variations, we investigated the reconstruction as a function of time on a yearly and daily basis, as shown in Fig. 8(a) and Fig. 8(b). We find small variations of the size of $2~{}\mathrm{g\,cm^{-2}}$ or $1~{}\mathrm{g\,cm^{-2}}$, respectively. We first calibrate the $X_{\mathrm{max}}$ predictions to remove the seasonal variation by fitting a sine function to the data and correcting it by using the time of year on an event-by-event basis. These predictions are also used to remove the diurnal variations by again fitting a sine wave and correcting events using Eq. 4. The dependencies after correction are depicted in Fig. 8(d) and Fig. 8(e).